can you please show me how to work this problem? 5[3(7-t)-4(8+2t)]-20=6[2(6+3t)-4] and 2+14x-9=7(2x+1)-14





Answer by stanbon(48535)

5[3(7-t)-4(8+2t)]-20=6[2(6+3t)-4]
5[21-3t-32-8t]-20=6[12+6t-4]
5[-11-11t]-20=6[8+6t]
-55-55t-20=48+36t
-123=-91t
t=123/91
-----------------------
2+14x-9=7(2x+1)-14
-7+14x=14x+7-14
-14=-14
This means x may have any value.
Cheers,
Stan H.